Gerber Clothing Inc. has designed a rain suit for outdoor enthusiasts that is about to be introduced on the market. A standard cost card has been prepared for the new suit, as follows:Gerber Clothing Inc. has designed a rain suit for outdoor enthusiasts that is about to be introduced on the market. A standarRequired: 1. Assume that the company uses the absorption approach to cost-plus pricing. a. Compute the markup that the companC-2. Compute the companys ROI for the year on the suits, using the ROI formula. (Do not round intermediate calculations.) RO
Answer #1

1.

a.

Total time available for the production of the new suit = 10,300 hours

Labour hour required per suit = 1 hour

Total number of suits could be manufactured = 10,300 / 1 = 10,300

Direct Materials ( 36 * 10,300) Direct Labour (40 * 10,300) Prime cost Add: Manufacturing Overhead ( 33 * 10,300) Manufacturi

Markup Percentage = (Required return + Sales & Administration) / Manufacturing Cost

= ( 106,000 + 623,755 ) / 1,122,700

= 65%

.

Note:

Under absorption costing system, the cost base for markup should be the manufacturing cost.

b.

Direct Materials Direct Labour Add: Manufacturing Overhead Cost of production/ per unit Add: Profit/Mark up Target Sales/ Sel

Target selling price = 179.85

c1.

Income statement Sales Less : Cost of goods sold Gross Margin Less : Selling, General & Administrative expenses 1,852,455 1,1

c2.

Company's ROI = 106,000 / 530,000 = 20%

.

2.

Mark up percentage for the total variable costing Sales units 10,300 Direct Materials 370,800 Direct Labor 412,000 Variable M